# RALL-E: Robust Codec Language Modeling with Chain-of-Thought Prompting for Text-to-Speech Synthesis

### v2

@@ 4.2 Main results
We first evaluate the overall performance of RALL-E on the full LibriSpeech test set with $1205$ utterances. The results are demonstrated in Table 2. It can be observed that RALL-E outperforms all other methods on WER. The reranked WER (WER-R) of RALL-E is even better than the WER of GT. Compared to the baseline VALL-E method, RALL-E obtains a $50\%$ relative improvement on WER and a $41\%$ relative improvement on WER-R, showing the superior robustness of the proposed method. This is also proved by the performance on the three error types, where RALL-E consistently reduces all error types from $2.8$/$1.5$/$1.3$ to $1.7$/$0.8$/$0.3$, respectively. In addition, the higher UTMOS score of RALL-E than the one of VALL-E indicates RALL-E can synthesize speech with better naturalness, showing the effectiveness of the proposed method on stabilizing speech prosody. Regarding the speaker similarity, both RALL-E and VALL-E show better performance than the previous two works which is possibly because of the larger training set used by us. However, the original VALL-E has a significantly higher SIM score ($0.58$) than other methods. One possible reason is that the speaker similarity of the original VALL-E is computed between the synthesized utterance and the prompt resynthesized by the codec instead of the GT prompt. We notice that VALL-T has slightly fewer insertion errors ($0.2$) than RALL-E ($0.3$), but we think this is reasonable and doesn’t necessarily mean RALL-E is worse than VALL-T since the test set of VALL-T only contains $500$ samples and fewer test samples usually result in a better WER, as suggested by the result of RALL-E ($904$) computed on the $904$ samples used by ELLA-V.
@@ 4.3 Ablation study
RALL-E $\mathbf{2.8}$ $\mathbf{4.00}$ $\mathbf{1.7}$ $0.8$ $\mathbf{0.3}$
w/o pitch $2.8$ $3.96$ $1.8$ $\mathbf{0.7}$ $0.3$
w/o window masking $2.9$ $3.84$ $1.8$ $0.8$ $0.3$
w/o duration-guided masking $3.4$ $3.88$ $2.0$ $0.9$ $0.5$
@@
We conduct ablation experiments to study the contributions of each component of RALL-E. Specifically, we study the following four settings: (1) w/o pitch that removes the pitch prompt; (2) w/o window masking where the window size $k$ is set to 00; (3) w/o duration-guided masking that uses normal unidirectional autoregressive attention masks; (4) w/o duration CoT prompting that removes duration from the CoT and model it separately. In the w/o duration CoT prompting setting we use an independent $8$-layer Transformer with $256$-dimensinal token embeddings and $8$ self-attention heads to predict duration, the masking strategy is still used based on the duration.
The results are shown in Table 4. First, the result of w/o pitch demonstrates that pitch has little contribution to the performance. While pitch prompting slightly improves speech naturalness and reduces mispronunciation, it causes more word omissions. Second, it can be observed from the result of w/o window masking that the model performance degrades when the window size is set to 00, which demonstrates the effectiveness of the window masking strategy. See Appendix A for a comprehensive study on the effects of window size $k$. Third, w/o duration-guided masking shows consistently worse performance than RALL-E on all metrics, showing the effectiveness of the proposed duration-guided masking strategy. Finally, w/o duration CoT prompting has the worst performance even though it uses the proposed duration-guided masking. By manually listening to the samples we find that the predicted duration of the independent Transformer does not influence the synthesized speech, which demonstrates the necessity of incorporating the duration with CoT prompting. In summary, each component of RALL-E contributes to the robustness improvements, and the CoT prompting is the most important component of RALL-E.
@@ 4.4 Evaluations on hard sentences
To further evaluate the robustness of RALL-E, we synthesize $50$ particularly hard sentences (see Appendix B for the transcripts of these sentences) with RALL-E and VALL-E. We manually evaluate the results since the WER computed on these sentences with a lot of numbers and symbols is imprecise. We decompose the possible errors into four types: mispronunciation, omission, repetition, and hallucination. Each utterance is synthesized $5$ times and the best one is selected. We count the frequency of all error types and compute the overall sentence error rate. Each error type is only counted once in an utterance. The results are shown in Table 1, in which we also add a powerful non-autoregressive TTS method NaturalSpeech2 (Shen et al., 2023) for reference. It can be seen that RALL-E substantially reduces the error rate from $68\%$ to $4\%$ with only $2$ mispronunciation errors. This performance approaches the one of non-autoregressive NaturalSpeech2 that has no error, which again demonstrates the effectiveness of RALL-E in improving the robustness of LLM-based TTS. Specifically, we observe that for extremely short sentences, e.g. a single alphabet “A”, VALL-E usually synthesizes words that don’t exist in the input sentence, causing a severe hallucination problem. Besides, if a word is repeated many times in a sentence, e.g. “$22$”, VALL-E easily makes mistakes by omitting or repeating the word. These problems together demonstrate that LLM-based TTS like VALL-E has low controllability on the duration of the synthesized speech and learns poor alignments between phonemes and speech tokens. In contrast, RALL-E first introduces prosody tokens as CoT prompting to improve duration controllability and then uses duration-guided masking to enhance the alignment, which effectively alleviates the aforementioned problems made by the language models. All in all, our RALL-E demonstrates superior robustness in all evaluations.

### v3

@@ 4.2 Main results
We first evaluate the overall performance of RALL-E on the full LibriSpeech test set with $1205$ utterances. The results are demonstrated in Table 2. It can be observed that RALL-E outperforms all other methods on WER. The reranked WER (WER-R) of RALL-E is even better than the WER of GT. Compared to the baseline VALL-E method, RALL-E obtains a $55\%$ relative improvement on WER and a $41\%$ relative improvement on WER-R, showing the superior robustness of the proposed method. This is also proved by the performance on the three error types, where RALL-E consistently reduces all error types from $2.8$/$1.5$/$1.3$ to $1.7$/$0.6$/$0.3$, respectively. In addition, the higher UTMOS score of RALL-E than the one of VALL-E indicates RALL-E can synthesize speech with better naturalness, showing the effectiveness of the proposed method on stabilizing speech prosody. Regarding the speaker similarity, both RALL-E and VALL-E show better performance than the previous two works which is possibly because of the larger training set used by us. However, the original VALL-E has a significantly higher SIM score ($0.58$) than other methods. One possible reason is that the speaker similarity of the original VALL-E is computed between the synthesized utterance and the prompt resynthesized by the codec instead of the GT prompt. We notice that VALL-T has slightly fewer insertion errors ($0.2$) than RALL-E ($0.3$), but we think this is reasonable and doesn’t necessarily mean RALL-E is worse than VALL-T since the test set of VALL-T only contains $500$ samples and fewer test samples usually result in a better WER, as suggested by the result of RALL-E ($912$) computed on the $912$ samples used by ELLA-V.
@@ 4.3 Ablation study
RALL-E $\mathbf{2.5}$ $\mathbf{4.00}$ $\mathbf{1.7}$ $\mathbf{0.5}$ $\mathbf{0.3}$
w/o pitch $2.6$ $3.96$ $1.8$ $0.5$ $0.3$
w/o window masking $2.7$ $3.84$ $1.8$ $0.6$ $0.3$
w/o duration-guided masking $3.2$ $3.88$ $2.0$ $0.8$ $0.5$
@@
We conduct ablation experiments to study the contributions of each component of RALL-E. Specifically, we study the following four settings: (1) w/o pitch that removes the pitch prompt; (2) w/o window masking where the window size $k$ is set to 00; (3) w/o duration-guided masking that uses normal unidirectional autoregressive attention masks; (4) w/o duration CoT prompting that removes duration from the CoT and model it separately. In the w/o duration CoT prompting setting we use an independent $8$-layer Transformer with $256$-dimensional token embeddings and $8$ self-attention heads to predict duration, the masking strategy is still used based on the duration.
The results are shown in Table 4. First, the result of w/o pitch demonstrates that the pitch token helps to reduce mispronunciation. Second, it can be observed from the result of w/o window masking that the model performance degrades when the window size is set to 00, which demonstrates the effectiveness of the window masking strategy. See Appendix A for a comprehensive study on the effects of window size $k$. Third, w/o duration-guided masking shows consistently worse performance than RALL-E on all metrics, showing the effectiveness of the proposed duration-guided masking strategy. Finally, w/o duration CoT prompting has the worst performance even though it uses the proposed duration-guided masking. By manually listening to the samples we find that the predicted duration of the independent Transformer does not influence the synthesized speech, which demonstrates the necessity of incorporating the duration with CoT prompting. In summary, each component of RALL-E contributes to the robustness improvements, and the CoT prompting is the most important component of RALL-E.
@@ 4.4 Evaluations on hard sentences
To further evaluate the robustness of RALL-E, we synthesize $50$ particularly hard sentences (see Appendix B for the transcripts of these sentences) with RALL-E and VALL-E. We manually evaluate the results since the WER computed on these sentences with a lot of numbers and symbols is imprecise. We decompose the possible errors into four types: mispronunciation, omission, repetition, and hallucination. Each utterance is synthesized $5$ times and the best one is selected. We count the frequency of all error types and compute the overall sentence error rate. Each error type is only counted once in an utterance. The results are shown in Table 1, in which we also add a powerful non-autoregressive TTS method NaturalSpeech2 [25] for reference. It can be seen that RALL-E substantially reduces the error rate from $68\%$ to $4\%$ with only $2$ mispronunciation errors. This performance approaches the one of non-autoregressive NaturalSpeech2 that has no error, which again demonstrates the effectiveness of RALL-E in improving the robustness of LLM-based TTS. Specifically, we observe that for extremely short sentences, e.g. a single alphabet “A”, VALL-E usually synthesizes words that don’t exist in the input sentence, causing a severe hallucination problem. Besides, if a word is repeated many times in a sentence, e.g. “$22$”, VALL-E easily makes mistakes by omitting or repeating the word. These problems together demonstrate that LLM-based TTS like VALL-E has low controllability on the duration of the synthesized speech and learns poor alignments between phonemes and speech tokens. In contrast, RALL-E first introduces prosody tokens as CoT prompting to improve duration controllability and then uses duration-guided masking to enhance the alignment, which effectively alleviates the aforementioned problems made by the language models. All in all, our RALL-E demonstrates superior robustness in all evaluations.